# Double and Triple Higgs Production to probe the Electroweak Phase Transition

### v1

@@ I Introduction
As the Higgs boson programme at the Large Hadron Collider (LHC) is entering a period of intense data gathering, rare processes become experimentally accessible and, hence, relevant. A particularly motivated process in this regard is the production of multiple Higgs bosons; di-Higgs boson production and associated analyses are cornerstones of the Beyond the Standard Model (BSM) physics programme at the LHC and beyond. Searches for Higgs pair production in gluon fusion $gg\to hh$ and weak boson fusion $pp\to hhjj$ are fairly well-established experimentally, cf. e.g. [1, 2]. With that in mind, both the theory and the experimental communities have started to increasingly look towards sensitivity to triple Higgs production. Perhaps a stretch goal in the SM context for the LHC even during its high-luminosity (HL) phase, triple Higgs production could indeed be a strategic target for future colliders [3, 4, 5, 6, 7, 8, 9, 10, 11]. Most targeted search strategies relying on a variety of optimisation methods and development, cf. e.g. [12, 13], will be the key to background mitigation and signal isolation of these final states.
@@ III Results and Discussion
How do these observations relate to the strength of the first-order phase transition? As a thermodynamic process, the EWPT is driven by the physics of the light degrees of freedom. For the concrete example of the R2HDM with relatively rigid coupling constraints of the Higgs bosons to other matter (in particular when considering the Higgs signal strength constraints inferred from LHC measurements), this is achieved by making the non-SM degrees of freedom more accessible via lighter spectra. In the opposite direction, when making the exotic states increasingly heavy by moving to the decoupling regime, we recover the electroweak cross-over of the SM Higgs sector. This is visible from Fig. 3, which distils the results of a scan of the strength of the phase transition $\xi_{p}$ at the percolation stage as a function of the mass of the neutral CP-even heavy Higgs boson $H$.
@@
Adding additional resonances to the spectrum, however, opens up significant parameter space for non-standard phenomenology. This can be observed in the cross-sections in the complex generalisation of the R2HDM, the C2HDM, along $hhh$ compared to $hh$, which is overlayed in Fig. 1. The cross section enhancements discussed above, i.e. $H\to hh$ and $H\to hhh$, can appear simultaneously when the two additional scalar degrees of freedom in the C2HDM are kinematically accessible in two- and three-body decays, respectively. The phenomenology of the $hh(h)$ final states is more involved. Due to the stringent limits from the measurements of the electric dipole moment, however, [80] (cf. the discussion of the impact on the C2HDM in [81, 82]) complex phases of the top- and bottom-Higgs interactions are quantitatively small as are the Higgs mixings that are absent in the R2HDM, so that no dramatic departure from the R2HDM paradigm is observed (or expected). The maximum amount of CP admixture observed for the SM-like scalar in the C2HDM is $\sim 9.5\%$ for $\xi_{p}>1$ and $\sim 10\%$ for $\xi_{p}<1$.
Similar to the C2HDM scenario, the distinctive resonance structures in the $M_{hh}$ and $M_{hhh}$ distributions then lead to enhancements of the triple Higgs production rate $\sim 20$ compared to di-Higgs production in the N2HDM (Fig. 4). Most notably, some of these parameter points do not show overly anomalous behaviour in the $hh$ modes; factors of $\lesssim 10$ enhancements will be extremely challenging at the HL-LHC, but represent an opportunity at, e.g., the FCC-hh. To this end, we note that all of our findings directly generalise to a potential FCC-hh, with reference cross sections of $\sigma_{hh}^{\text{SM}}\approx 800~{}\text{fb}$ [83] and $\sigma_{hhh}^{\text{SM}}\approx 2.9~{}\text{fb}$ [8].

### v2

@@ I Introduction
As the Higgs boson programme at the Large Hadron Collider (LHC) is entering a period of intense data gathering, rare processes become experimentally accessible and, hence, relevant. A particularly motivated process in this regard is the production of multiple Higgs bosons; di-Higgs boson production and associated analyses are cornerstones of the Beyond the Standard Model (BSM) physics programme at the LHC and beyond. Searches for Higgs pair production in gluon fusion $gg\to hh$ and weak boson fusion $pp\to hhjj$ are fairly well-established experimentally, cf. e.g. [1, 2]. With that in mind, both the theory and the experimental communities have started to increasingly look towards sensitivity to triple Higgs production. Perhaps a stretch goal in the SM context for the LHC even during its high-luminosity (HL) phase, triple Higgs production could indeed be a strategic target for future colliders [3, 4, 5, 6, 7, 8, 9, 10, 11, 12]. Most targeted search strategies relying on a variety of optimisation methods and development, cf. e.g. [13, 14], will be the key to background mitigation and signal isolation of these final states.
@@ III Results and Discussion
How do these observations relate to the strength of the first-order phase transition? As a thermodynamic process, the EWPT is driven by the physics of the light degrees of freedom. For the concrete example of the R2HDM with relatively rigid coupling constraints of the Higgs bosons to other matter (in particular when considering the Higgs signal strength constraints inferred from LHC measurements), this is achieved by making the non-SM degrees of freedom more accessible via lighter spectra. This is visible from Fig. 3, which distils the results of a scan of the strength of the phase transition $\xi_{p}$ at the percolation stage as a function of the mass of the neutral CP-even heavy Higgs boson $H$.
@@
Adding additional resonances to the spectrum, however, opens up significant parameter space for non-standard phenomenology. This can be observed in the cross-sections in the complex generalisation of the R2HDM, the C2HDM, along $hhh$ compared to $hh$, which is overlayed in Fig. 1. The cross section enhancements discussed above, i.e. $H\to hh$ and $H\to hhh$, can appear simultaneously when the two additional scalar degrees of freedom in the C2HDM are kinematically accessible in two- and three-body decays, respectively. The phenomenology of the $hh(h)$ final states is more involved. Due to the stringent limits from the measurements of the electric dipole moment, however, [83] (cf. the discussion of the impact on the C2HDM in [84, 85]) complex phases of the top- and bottom-Higgs interactions are quantitatively small as are the Higgs mixings that are absent in the R2HDM, so that no dramatic departure from the R2HDM paradigm is observed (or expected). This can also be visualised looking at the amount of CP admixture, which is a measure of the SM-like scalar-pseudoscalar mixing. For the Type-I C2HDM, following Refs. [86, 84], the pseudoscalar admixture is defined as,
$\Psi_{i}^{\text{C2HDM}}=(R_{i3})^{2},$ (15)
where $R$ is the orthogonal rotation matrix that diagonalises the neutral mass-mixing matrix. The maximum amount of CP admixture observed for the SM-like scalar in the C2HDM is $\sim 9.5\%$ for $\xi_{p}>1$ and $\sim 10\%$ for $\xi_{p}<1$.
Similar to the C2HDM scenario, the distinctive resonance structures in the $M_{hh}$ and $M_{hhh}$ distributions then lead to enhancements of the triple Higgs production rate $\sim 20$ compared to di-Higgs production in the N2HDM (Fig. 4). Most notably, some of these parameter points do not show overly anomalous behaviour in the $hh$ modes. This can be attributed to the fact that our scan is statistically limited by the number of points with the heavier exotics having a mass close to the $2m_{h}$ threshold, where we observe the most enhancements in double Higgs production. Factors of $\lesssim 10$ enhancements in triple Higgs will be extremely challenging at the HL-LHC but represent an opportunity at, e.g., the FCC-hh. To this end, we note that all of our findings directly generalise to a potential FCC-hh, with reference cross sections of $\sigma_{hh}^{\text{SM}}\approx 800~{}\text{fb}$ [87] and $\sigma_{hhh}^{\text{SM}}\approx 2.9~{}\text{fb}$ [8].
Our findings are based on a leading-order exploration of the electroweak sector. Whilst QCD corrections are known to be relatively insensitive to the mass scales in multi-Higgs production (which is, in fact, exploited in most precision calculations there), weak radiative corrections might sculpt our findings quantitatively, see [88, 89, 90, 80] for the impact on the 125 GeV self-coupling. These modifications are only part of the full (resonant) amplitude and warrant further investigation. Our observation related to the lightness of the degrees of freedom should be relatively robust against higher-order corrections.